e re he ou ve a complehe the hip? s the Pn d he
Answer #1

Let A: You go by mule

B: You go by walking

C : You go by flying

D: you complete ths trip

Given: P( A)= 0.5 , P(B)= 0.1, P(C)=0.4, P(D/A)=0.10, P( D/C)=0.9, P(D/B)= 0.01

aue p(BAD) 9 b) ОIo x Oi5-t ο.nlXQLl+ O.AX Oay P (5) P (5) ー-PIA)-p(DIP).PIA) - ㄧ一ㄧ rib):-PLD)一 1-0.91) = 0,539 0.531